POYO:

If y varies directly with x and x=-16 when y = 28,what is the direct variation equation.

jhonyy9:

so i will try explain it to you directly variation mean a variation in the same way for example x = -16 and y = 28 when will be doubled x so than x = -32 in this way y = 28 +28 = 56 doubled the same like x hope you understand me

POYO:

yeah sorta but yeah

jhonyy9:

so in this way after this above my example can you write the direct variation equation ?

jhonyy9:

how will be this equation ?

jhonyy9:

may be this right x = y 2x = 2y 3x = 3y ??? or maybe x = 1/y or 1/x = y ?

jhonyy9:

please read my above wrote example and than you understand so will get the right correct equation easy

POYO:

well i have some answer that i think are right but its 4

jhonyy9:

4 equations ?

jhonyy9:

you need just 1

POYO:

i will show you

jhonyy9:

ok

POYO:

y=-7/4x y=12x y=7/4x 28=-16k

POYO:

here

POYO:

i know that one of them are right but i dont know

jhonyy9:

x=-16 when y = 28 direct variation mean the same variation of every terms -16x = 28y how can you separate the y in a part - side ?

POYO:

so its 28=-16

jhonyy9:

do you see this true ?

POYO:

yeah

POYO:

i think

jhonyy9:

for you is indifferent than you have 28 dollars or you need give to somebody 16 ?

jhonyy9:

x=-16 when y = 28 28y = -16x so to separate the y in on a side you divide both sides by 28 so in this way what will get ?

POYO:

ok thx

POYO:

i am going to close this now

jhonyy9:

really ? why ?

POYO:

you helped me and now i get it

jhonyy9:

ok than what is the right answer equation ?